# Multi-color x-ray free-electron laser generation using optical klystron for multi-frame diffraction imaging

### v2

@@ I Introduction
Various multi-color XFEL schemes have been proposed by manipulating electron beam properties, including twin bunches [29], two-bucket bunches [11], and the use of different energy parts within one bunch [44], double-slot foil [45], sextupole [13], nonlinear electron compression [28] and laser emittance spoiler [48]. Meanwhile, artificially changing the undulator parameter can straightforwardly produce tunable multi-color XFEL pulses [26, 20, 10]. However, when the whole electron bunch is used to generate multiple colors sequentially, the intensity of downstream pulses is limited by the energy spread induced by upstream lasing. The fresh-slice technique circumvents this limitation by selectively utilizing unmodulated portions of the beam for independent amplification in separate undulators, thereby enabling high-power multi-color operation [27, 43, 19, 7, 39].
@@ II Principles
When the bunched electron beam passes through the sub-undulator $U_{A,N-1}$, it undergoes N2N-2 stages of OK-enhanced SASE FEL amplification, specifically referred to as the seeded FEL amplification. During this process, the energy modulation amplitude induced by the sub-undulator $U_{A,N-1}$ before the last dispersive chicane $R_{56}^{N-1}$ becomes larger than the intrinsic energy spread. As the energy modulation amplitude increases, the optimal value of $R_{56}^{N-1}$ decreases due to the dominance of the Bessel function governing the bunching process. Specifically, a larger energy modulation amplitude requires a smaller value $R_{56}^{N-1}$ to maximize the bunching factor, as shown in Fig. 3. This optimization ultimately enhances FEL radiation in the subsequent sub-undulator $U_{A,N}$. The time delay of delay chicane 1 is twice its $R_{56}$, which is sufficient to smear out the microbunching generated by the previous undulator module $A_{N}$. For instance, when $R_{56}=600\ \mu m$, the bunching factor is nearly zero, as given by Eq. (3). Similarly, the multi-stage OK configuration involves N1N-1 small dispersive chicanes for each wavelength, including $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$.
@@
In the proposed scheme, adjusting the time delay of the large chicane controls the inter-pulse time separation between adjacent X-ray pulses to set the frame interval. Tuning the $K$ parameter of sub-undulators regulates both the wavelength separation and specific X-ray wavelengths. The energy of each X-ray pulse can be effectively controlled by jointly tuning the number of sub-undulators and the parameters $R_{56}$ of dispersive chicanes, where $R_{56}$ is optimized by the differential evolution algorithm – a stochastic population-based algorithm employing vector differential mutation, binomial crossover, and competitive selection mechanisms [53]. Theoretically, this scheme enables the generation of multi-color X-ray pulses for multi-frame diffraction imaging. However, limitations in adjustable frame numbers exist due to electron beam energy spread growth and the dynamic range constraints of the X-ray CCD, which are further influenced by its physical size.
@@ III Start-to-end XFEL simulation
Under fully optimized $R_{56}$ values for all dispersive chicanes, the undulator modules $A_{5}$, $B_{6}$, $C_{6}$, $D_{6}$ (with five sub-undulators in $A_{5}$ and six sub-undulators in each of $B_{6}$, $C_{6}$ and $D_{6}$) are operated with undulator parameters $K_{1}$ = 2.00, $K_{2}$ = 2.21, $K_{3}$ = 2.39 and $K_{4}$ = 2.56, as illustrated in Fig. 5(c). The electron beam propagating through these undulator modules generates radiation at the resonant wavelengths of 1.8, 2.1, 2.4, and 2.7 nm, with adjacent wavelengths separated by 0.3 nm, shown in Fig. 5(a). These wavelengths correspond to photon energies of 688.80, 590.40, 516.60, and 459.20 eV, respectively. Independent generation of each X-ray pulse is enabled by their time separation via delay chicanes, resulting in inter-pulse temporal separations of delay = 1 ps, evident in the FEL power profile (see Fig. 5(b)). All four-color X-ray pulses achieve peak power levels on the order of hundreds of megawatts. To realize this multi-color output within a single pass, the same electron bunch is reused across multiple undulator modules. However, the FEL radiation in the upstream sections induces significant energy spread in the high-current core of the bunch, thereby degrading its FEL gain for downstream wavelengths. Consequently, lasing at subsequent colors is sustained primarily by the low-energy-spread electrons in the bunch head and tail. This longitudinal partitioning of lasing regions leads to temporal pulse splitting in the output. As demonstrated in the following sections, a precisely balanced pulse energy configuration across colors effectively suppresses this splitting phenomenon, thereby improving temporal structure and synchronization.
@@
As the beam energy spread increases along the undulator modules in Fig. 6(b), the condition $\sigma_{\delta}\ll\rho$, to which the OK enhancement is highly sensitive [4, 14, 38], is progressively violated, leading to declining pulse energies for the different X-ray wavelengths: 30.4 $\mu J$ (1.8 nm), 25.6 $\mu J$ (2.1 nm), 1.9 $\mu J$ (2.4 nm), and 2.9 $\mu J$ (2.7 nm). In contrast, when OK chicanes are removed, the corresponding pulse energies drop to 0.2, 4.5, 1.7, and 5.5 $\mu J$. The multi-stage OK enhancement is most effective for the first three pulses. For the fourth pulse, the case with OK does not yield significant FEL improvement due to its larger initial energy spread, compared to the scenario without OK (dot-dashed line), which exhibits a more gradual energy spread growth and thus a relatively smaller initial energy spread. However, when the first three pulses are enhanced by OK while the fourth pulse reverts to normal SASE (gray dashed line), the output performance deteriorates significantly. This indicates that the OK mechanism still plays a crucial role in accelerating the FEL gain process across different wavelengths. As a result, high-power four-color XFEL pulses can be generated within shorter undulator modules.
@@ V Conclusion and discussion
A novel single-shot multi-frame X-ray diffraction imaging scheme is proposed by theoretical analysis and start-to-end simulation. The scheme employs the whole electron bunch in a multi-stage OK-SASE configuration with split undulators to generate high-power, four-color XFEL pulses featuring independently tunable wavelength separations and controllable inter-pulse delays. Delay chicanes placed in the gaps between undulator modules operating at different resonant wavelengths enable time delays on the order of one hundred femtoseconds, extendable into the picosecond regime. The feasibility of the imaging process is confirmed by X-ray transport simulations, which demonstrate that the resulting diffraction signals from distinct probe pulses can be spatially resolved and recorded simultaneously, thereby validating the scheme’s capability for capturing ultrafast dynamics in a single shot.
@@
Furthermore, the capability to flexibly generate wavelengths of interest, such as 1.8, 2.1, 2.4, and 2.7 nm, with a tunable separation (e.g., 0.3 nm) plays an important role in optimizing both the contrast and spatial resolution of diffraction imaging. This flexibility is particularly advantageous for probing elemental absorption edges, for example, the oxygen K-edge at 529 eV. Moreover, extending the scheme to other soft X-ray regions, such as around 6 nm covering the carbon K-edge at 284 eV, further broadens its applicability to a wide range of scientific problems. Moreover, the pulse energy of each color can be effectively controlled through precise adjustment of the $R_{56}$ values of the dispersive chicanes. This enables the delivery of sufficient photon flux to improve the signal-to-noise ratio and imaging fidelity, while simultaneously mitigating radiation damage to samples, which is especially beneficial for studies of fragile specimens. Since all pulses are generated from a single electron bunch, independent control of their temporal profiles is inherently limited, and residual energy modulation may induce pulse splitting in subsequent undulator modules. This effect may be mitigated by reducing the radiation intensity in upstream stages while gradually increasing it in downstream stages, or by enlarging the wavelength separation between adjacent pulses. In addition, the residual beam energy chirp of about 0.22% remains smaller than the FEL gain bandwidth of 0.34%, so no significant spectral broadening is observed. The implementation of a dechirper would allow more precise manipulation of the longitudinal phase space of the electron beam, offering further flexibility in tailoring pulse characteristics.
@@
The multi-stage OK configuration can be extended to other multi-color approaches, such as fresh-slice operation, where it can further enhance FEL gain and enable higher peak power or shorter pulses. In particular, under low-charge operation, where fresh-slice manipulation becomes challenging, the full-beam scheme combined with OK may still provide a practical route to two-color FEL generation. Nevertheless, the proposed scheme is expected to provide a better foundation for designing sample-specific light sources tailored to dynamic imaging experiments for XFEL users, with its multi-color X-ray pulses enabling applications such as multi-color pump-probe experiments, nonlinear X-ray spectroscopy, multi-wavelength anomalous diffraction, and other multi-wavelength techniques. Collectively, these capabilities provide a robust framework for exploring microscopic mechanisms in ultrafast science.

### v3

@@ I Introduction
Various multi-color XFEL schemes have been proposed by manipulating electron beam properties, including twin bunches [29], two-bucket bunches [11], and the use of different energy parts within one bunch [44], double-slot foil [45], sextupole [13], nonlinear electron compression [28] and laser emittance spoiler [48]. Meanwhile, artificially changing the undulator parameter can straightforwardly produce tunable multi-color XFEL pulses [26, 20, 10]. However, when the whole electron bunch is used to generate multiple colors sequentially, the intensity of downstream pulses is limited by the energy spread induced by upstream lasing. The fresh-slice technique circumvents this limitation by selectively utilizing unmodulated portions of the beam for independent amplification in separate undulators, thereby enabling high-power multi-color operation [27, 43, 19, 7, 39, 49].
@@ II Principles
When the bunched electron beam passes through the sub-undulator $U_{A,N-1}$, it undergoes N2N-2 stages of OK-enhanced SASE FEL amplification, specifically referred to as the seeded FEL amplification. During this process, the energy modulation amplitude induced by the sub-undulator $U_{A,N-1}$ before the last dispersive chicane $R_{56}^{N-1}$ becomes larger than the intrinsic energy spread. As the energy modulation amplitude increases, the optimal value of $R_{56}^{N-1}$ decreases due to the dominance of the Bessel function governing the bunching process. Specifically, a larger energy modulation amplitude requires a smaller value $R_{56}^{N-1}$ to maximize the bunching factor, as shown in Fig. 3. This optimization ultimately enhances FEL radiation in the subsequent sub-undulator $U_{A,N}$.
@@
Additionally, the OK enhancement is also highly sensitive to the initial slice energy spread [4, 14, 38]. As shown in Fig. 4, the pulse energy decreases rapidly as the initial slice energy spread increases, indicating clear performance degradation. For each point, the corresponding $R_{56}$ values were re-optimized using the differential evolution algorithm [54] to maximize the pulse energy at $\lambda_{1}$.
Figure 4: Dependence of the pulse energy at 1.8 nm on the initial slice energy spread. The beam energy and peak current are fixed at 4.5 GeV and 800 A, respectively.
The time delay of delay chicane 1 is twice its $R_{56}$, which is sufficient to smear out the microbunching generated by the previous undulator module $A_{N}$. For instance, when $R_{56}=600\ \mu m$, the bunching factor is nearly zero, as given by Eq. (3). Similarly, the multi-stage OK configuration involves N1N-1 small dispersive chicanes for each wavelength, including $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$.
@@
In the proposed scheme, adjusting the time delay of the large chicane controls the inter-pulse time separation between adjacent X-ray pulses to set the frame interval. Tuning the $K$ parameter of sub-undulators regulates both the wavelength separation and specific X-ray wavelengths. The energy of each X-ray pulse can be effectively controlled by jointly tuning the number of sub-undulators and the optimized $R_{56}$ values of dispersive chicanes. Theoretically, this scheme enables the generation of multi-color X-ray pulses for multi-frame diffraction imaging. However, limitations in adjustable frame numbers exist due to electron beam energy spread growth and the dynamic range constraints of the X-ray CCD, which are further influenced by its physical size.
@@ III Start-to-end XFEL simulation
Under fully optimized $R_{56}$ values for all dispersive chicanes, the undulator modules $A_{5}$, $B_{6}$, $C_{6}$, $D_{6}$ (with five sub-undulators in $A_{5}$ and six sub-undulators in each of $B_{6}$, $C_{6}$ and $D_{6}$) are operated with undulator parameters $K_{1}$ = 2.00, $K_{2}$ = 2.21, $K_{3}$ = 2.39 and $K_{4}$ = 2.56, as illustrated in Fig. 6(c). The electron beam propagating through these undulator modules generates radiation at the resonant wavelengths of 1.8, 2.1, 2.4, and 2.7 nm, with adjacent wavelengths separated by 0.3 nm, shown in Fig. 6(a). These wavelengths correspond to photon energies of 688.80, 590.40, 516.60, and 459.20 eV, respectively. The four-color X-ray pulses achieve peak power levels on the order of hundreds of megawatts, as illustrated in Fig. 6(b). Independent generation of each X-ray pulse is enabled by delay chicanes, which provide a temporal separation of 1 ps between neighboring pulses, corresponding to the maximum delay used to clearly separate the four pulses. The minimum useful delay is constrained by the XFEL pulse length, below which neighboring pulses would overlap and become difficult to distinguish. This inter-pulse delay defines the accessible temporal sampling points between successive pulses, but it does not directly determine the temporal resolution. Instead, the temporal resolution is mainly limited by the XFEL pulse length and is on the order of 100 fs for the present case, consistent with the values listed in Table 2. To realize this multi-color output within a single pass, the same electron bunch is reused across multiple undulator modules. However, the FEL radiation in the upstream sections induces significant energy spread in the high-current core of the bunch, thereby degrading its FEL gain for downstream wavelengths. Consequently, lasing at subsequent colors is sustained primarily by the low-energy-spread electrons in the bunch head and tail. This longitudinal partitioning of lasing regions leads to temporal pulse splitting in the output. As demonstrated in the following sections, a precisely balanced pulse energy configuration across colors effectively suppresses this splitting phenomenon, thereby improving temporal structure and synchronization.
@@
As the beam energy spread increases along the undulator modules in Fig. 7(b), the condition $\sigma_{\delta}\ll\rho$ is progressively violated, leading to declining pulse energies for the different X-ray wavelengths: 30.4 $\mu J$ (1.8 nm), 25.6 $\mu J$ (2.1 nm), 1.9 $\mu J$ (2.4 nm), and 2.9 $\mu J$ (2.7 nm). This behavior is consistent with the idealized sensitivity analysis shown in Fig. 4. In contrast, when OK chicanes are removed, the corresponding pulse energies drop to 0.2, 4.9, 1.7, and 5.5 $\mu J$. The multi-stage OK enhancement is most effective for the first three pulses. For the fourth pulse, the case with OK does not yield significant FEL improvement due to its larger initial energy spread, compared to the scenario without OK (dot-dashed line), which exhibits a more gradual energy spread growth and thus a relatively smaller initial energy spread. However, when the first three pulses are enhanced by OK while the fourth pulse reverts to normal SASE (gray dashed line), the output performance deteriorates significantly. This indicates that the OK mechanism still plays a crucial role in accelerating the FEL gain process across different wavelengths, thereby enabling four-color XFEL pulse generation within shorter undulator modules. The corresponding output parameters, including peak power, pulse length, and pulse energy, are summarized in Table 2.
Table 2: Main output parameters of the four XFEL pulses. Values in each entry are given as OK-SASE / SASE.
First pulse Second pulse Third pulse Fourth pulse
Peak power (GW) 1.72/0.02 2.55/0.42 0.53/0.14 0.42/0.75
Pulse length (fs) 31.6/44.7 66.5/39.0 71.7/53.1 70.2/48.1
[…]
@@ V Conclusion and discussion
A novel single-shot multi-frame X-ray diffraction imaging scheme is proposed by theoretical analysis and start-to-end simulation. The scheme employs the whole electron bunch in a multi-stage OK-SASE configuration with split undulators to generate four-color XFEL pulses featuring independently tunable wavelength separations and controllable inter-pulse delays, while reducing the required undulator length. Delay chicanes placed in the gaps between undulator modules operating at different resonant wavelengths enable time delays on the order of one hundred femtoseconds, extendable into the picosecond regime. The feasibility of the imaging process is confirmed by X-ray transport simulations, which demonstrate that the resulting diffraction signals from distinct probe pulses can be spatially resolved and recorded simultaneously, thereby validating the scheme’s capability for capturing ultrafast dynamics in a single shot.
@@
Furthermore, the capability to flexibly generate wavelengths of interest, such as 1.8, 2.1, 2.4, and 2.7 nm, with a tunable separation (e.g., 0.3 nm) plays an important role in optimizing both the contrast and spatial resolution of diffraction imaging. This flexibility is particularly advantageous for probing elemental absorption edges, for example, the oxygen K-edge at 529 eV. Moreover, extending the scheme to other soft X-ray regions, such as around 6 nm covering the carbon K-edge at 284 eV, further broadens its applicability to a wide range of scientific problems. In addition, the achievable temporal resolution is on the order of 100 fs, mainly limited by the XFEL pulse length, while this pulse length can be adjusted through accelerator operating modes. Under more advanced short-pulse operating conditions, it could be reduced to several tens of femtoseconds or even approach 10 fs level, thereby further improving the temporal resolution. The pulse energy of each color can be effectively controlled by precisely adjusting the $R_{56}$ values of the dispersive chicanes. Balanced pulse energy is about 10 $\mu J$ per color, which remains very low compared with standard single-color operation at existing XFEL facilities. Despite this limitation, such tunability may still be useful for distributing the available photon flux among different colors according to experimental needs and for reducing radiation damage to fragile samples. Since all pulses are generated from a single electron bunch, independent control of their temporal profiles is inherently limited, and residual energy modulation may induce pulse splitting in subsequent undulator modules. This effect may be mitigated by reducing the radiation intensity in upstream stages while gradually increasing it in downstream stages, or by enlarging the wavelength separation between adjacent pulses. In addition, the residual beam energy chirp of about 0.22% remains smaller than the FEL gain bandwidth of 0.34%, so no significant spectral broadening is observed. The implementation of a dechirper would allow more precise manipulation of the longitudinal phase space of the electron beam, offering further flexibility in tailoring pulse characteristics.
@@
The multi-stage OK configuration can also be extended to other multi-color approaches, such as fresh-slice operation, where it can further enhance FEL gain and enable higher peak power or shorter pulses. In particular, under low-charge operation, where fresh-slice manipulation becomes challenging, the full-beam scheme combined with OK may still provide a practical route to two-color FEL generation. Overall, the proposed scheme is expected to provide a better foundation for designing sample-specific light sources tailored to dynamic imaging experiments for XFEL users, with its multi-color X-ray pulses enabling applications such as multi-color pump-probe experiments, nonlinear X-ray spectroscopy, multi-wavelength anomalous diffraction, and other multi-wavelength techniques. Collectively, these capabilities provide a robust framework for exploring microscopic mechanisms in ultrafast science.